OpenStudy (liz_beauty45):

Evaluate for x = 2. (12x + 8) 4 A) 4 B) 5 C) 6 D) 8

OpenStudy (liz_beauty45):

plz plz plz plz plz help me

OpenStudy (fortytherapper):

\[(12)(x) +8 =?\] Replace x with 2, since x = 2 \[(12)(2) + 8 = ?\] There is no answer though, is there a typo anywhere? Could it be that 4 below your question?

OpenStudy (liz_beauty45):

14 and 8 is 22?

OpenStudy (sleepyjess):

Hello ^_^ Could you possibly type out the equation using the equation editor?

OpenStudy (fortytherapper):

The 12 and 2 are being multiplied. You would need PEMDAS. First multiply and then add

OpenStudy (fortytherapper):

You know what; is that 4 being divided?

OpenStudy (liz_beauty45):

oh so 12 timez 2 is 24 + 8 its 32?????

OpenStudy (fortytherapper):

Correct

OpenStudy (liz_beauty45):

now what

OpenStudy (fortytherapper):

Is it this; \[\frac{ (12x +2 }{ 4}\]

OpenStudy (liz_beauty45):

thankyou so much u are a big help

OpenStudy (fortytherapper):

32 isn't an answer choice though

OpenStudy (liz_beauty45):

yeah i noticed that

OpenStudy (fortytherapper):

Or is all that being divided by 4 \[\frac{ (12(2)+2) }{ 4 }\]

OpenStudy (fortytherapper):

+8 I mean

OpenStudy (liz_beauty45):

OH OKAY but wait a sec the answer would be 4 right?